OpenStudy (anonymous):

(1/2 - 1/4i) (5/6 + 5/6i) Help me Multiply. Please!

OpenStudy (espex):

Multiply like you would two factors, then add the real parts and add the imaginary parts.

OpenStudy (anonymous):

This is so hard for me. Can you please go through the steps and help me with the answer?

OpenStudy (anonymous):

@eSpeX Can you help me with this? This is really hard for me.

OpenStudy (espex):

Let's confirm, this is your equation, yes? \[(\frac{1}{2}-\frac{1}{4}i)(\frac{5}{6}+\frac{5}{6}i)\]

OpenStudy (anonymous):

Yes!

OpenStudy (espex):

You then foil the two terms like you would (x+2)(x-4)

OpenStudy (anonymous):

1/2 * 5/6 ?

OpenStudy (espex):

Yes

OpenStudy (anonymous):

15/6 ?

OpenStudy (espex):

5/12

OpenStudy (espex):

You multiply the denominators and the numerators.

OpenStudy (anonymous):

Oh! Alright, so far 5/12 ?

OpenStudy (anonymous):

Then would it be 5/12 again?

OpenStudy (espex):

Yes, but you have an i as well so it would be 5/12 i

OpenStudy (anonymous):

Would -1/4 i * 5/6 be 60/24i ?

OpenStudy (espex):

\[\frac{1*5}{4*6}i\]

OpenStudy (espex):

negative of course.

OpenStudy (anonymous):

5/24 i?

OpenStudy (espex):

Yes

OpenStudy (anonymous):

So far 5/12 + 5/12i - 5/24i + 5/24 i^2

OpenStudy (espex):

Exactly, now recall what 'i' stands for, it is the sqrt of -1, so i^2 is?

OpenStudy (anonymous):

5/24(-1) = -5/24?

OpenStudy (espex):

Precisely.

OpenStudy (anonymous):

So would the answer be 5/12 + 5/12i +5/24i - 5/24 ?

OpenStudy (espex):

You must now combine like terms.

OpenStudy (anonymous):

5/12 - 5/24 be 5/24?

OpenStudy (espex):

No, you would need to give them the same denominator to be able to subtract them. Also, it seems we dropped a sign somewhere, I think it was when you did -1/4i*5/6i, it would have given you -5/24(-1) to end up with 5/24.

OpenStudy (anonymous):

So the same Denominator would be 24?

OpenStudy (espex):

So what would you need to do to 5/12 to get the same denominator? Yes, 24.

OpenStudy (anonymous):

Multiply by the GCF?

OpenStudy (espex):

A common factor yes.

OpenStudy (anonymous):

15/24 ?

OpenStudy (espex):

Is that your conversion or final answer?

OpenStudy (anonymous):

My conversion

OpenStudy (espex):

How did you get it?

OpenStudy (anonymous):

multiply and adding?

OpenStudy (espex):

Okay, so you are saying it is your final answer, because your conversion of 5/12 was 10/24 and you added them together to get 15/24.

OpenStudy (anonymous):

Yes.

OpenStudy (espex):

Okay, now you will do the same thing to your imaginary parts. \[\frac{5}{12}i-\frac{5}{24}i\]

OpenStudy (anonymous):

5/24 i ?

OpenStudy (espex):

Good. So what is your final result?

OpenStudy (anonymous):

15/24 - 5/24 i ?

OpenStudy (espex):

And can you reduce any of those terms?

OpenStudy (anonymous):

no

OpenStudy (espex):

What about 15/24, don't they have a common factor?

OpenStudy (anonymous):

5/8

OpenStudy (espex):

Perfect. Excellent work.

OpenStudy (anonymous):

So the final answer is: 5/8 - 5/24 i ?

OpenStudy (espex):

Indeed it is.

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (espex):

You did all the heavy lifting, but you are certainly welcome.